OpenStudy (anonymous):

simplyfy (x-9)^3/5=8 how do you solve the power to 3/5?

OpenStudy (watchmath):

Raise both sides to the 5/3 power (x-9)=8^(5/3)=32 x-9=32 x=41

OpenStudy (anonymous):

i dont understand still

OpenStudy (anonymous):

\[\sqrt[5]{(x-9)^2} = (x-9)^{3/5}\]

OpenStudy (anonymous):

sorry, should be \[\sqrt[5]{(x-9)^{3}}\]

OpenStudy (anonymous):

Now\[[\sqrt[5]{(x-9)^{3}}]^{5} = [8]^5\] \[(x-9)^3 = 8^5\] \[\sqrt[3]{(x-9)^3} = \sqrt[3]{[8^5]}\] \[x-9 = 8^{5/3}\]

OpenStudy (anonymous):

\[8^{5/3} = \sqrt[3]{32768} = 32\]

OpenStudy (anonymous):

so x-9 = 32 add 9 both sides x-9+9=32+9 x=41 hope it helps!

OpenStudy (anonymous):

lets try this. rather than raising to the fifth power first, take the cube root of 8 first. the cube root of 8 is 2, and 2 to the fifth power is 32. much easier that way.